Volume occupied by molecules of one mole of a gas at NTP , each having a radius of 1Angstorm is

Grade:11

4 Answers

Susmita
5 years ago
At NTP 1 mole of a gas occupies a volume of  22.4 litre.
Also at NTP number of gas molecules in 1 mole of a gas is NA( Avogadro number 6.023×1023).
So volume occupied by a single molecule is
22.4/Nlitre.
Susmita
5 years ago
As the question says volume occupied by molecules of 1mole gas at NTP,it will simply be 22.4 lit.You do not have to calculate anything.
Chinmay
5 years ago
I found this ans.but how is it? 
Volume of 1 mole of molecules=N×(4v)
(b=4v)
=(6.02×10^23)(4×4/3 πr^3)
=(6.02×10^23)(4×4/3× 22/7 )×(10^-8cm)^3
=10.09 cm^3 or ml
Susmita
5 years ago
Why are you multiplying by 4?
The volume of a single molecule is (4/3)×\pi×r3
If you multiply this with 6.023×1023 you will get almost  25 litre which is close to 22.4 litre,